OpenStudy (anonymous):

(x)/(1-x) = (1-x)/(1) find sin of theta and cos of theta (x)/(1-x) = (1-x)/(1) find sin of theta and cos of theta @Mathematics

OpenStudy (anonymous):

\[find \sin \theta and \cos \theta\]

myininaya (myininaya):

i don't get it

OpenStudy (anonymous):